OpenStudy (anonymous):

Use the cross-product method to find out whether the pair of fractions is equivalent. 2/9, 4/16

OpenStudy (anonymous):

2*16 = 4 *9 32 = 32 cool

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

\[\frac{2}{9} \times \frac{4}{16}\] you multiply across the diagonals. the 2 matches with the 16 and the 9 matches with the 4.

OpenStudy (anonymous):

okay thanks so much i like it a lot better when people can help explain things to me , what about this one? divide using long divison 8 divided by 24,324

OpenStudy (anonymous):

=8

OpenStudy (anonymous):

it's hard to explain long division without talking through it, honestly =/ there's a nice wikipedia though. http://en.wikipedia.org/wiki/Long_division